# A practical identifiability criterion leveraging weak-form parameter estimation

### v4

@@ 3.2 Weak form Estimation of Nonlinear Dynamics (WENDy)
As presented in [BortzMessengerDukic2023BullMathBiol], the WENDy algorithm addresses the Errors-In-Variable nature of the linearized problem using iterative reweighting to approximate the correct underlying covariance structure [^8] The approximated covariance can also be used to calculate parameter confidence intervals. For a derivation of the parameter covariance approximation, we direct readers to [BortzMessengerDukic2023BullMathBiol].. This reweighting is dependent on the assumed noise structure of the observed data. For Gaussian additive noise, we apply the covariance correction using Algorithm 2 in [BortzMessengerDukic2023BullMathBiol]. Additionally, it has been shown in [RummelMessengerBeckerEtAl2025arXiv250208881] that, through a strategic choice of transformations to the observations $\mathbf{y}$, it is possible to approximate the covariance structure for systems where observational noise is multiplicative and lognormal (i.e. $\mathbf{y}=\mathbf{y}^{\star}\varepsilon$ where $\log(\varepsilon(\mathbf{t}))\sim\mathcal{N}(0,\sigma^{2})$).
@@ 5 Conclusion
While global practical identifiability analysis is an emerging area of research [BholaDuraisamy2023SciRep, DobreBastogneRichard2010IFACProceedingsVolumes, ElWajehJungBongartzEtAl2022BullMathBiol] our $(e,q)$ criterion, like profile likelihood analysis and the FIM method, the $(e,q)$ method is local. This implies that, following best modeling practices, one may need to test a range of values in the parameter space rather than near a single parameterization of a system. For example, as we demonstrate in Appendix E, depending on the dynamics of the blood concentration, the parameter $w_{3}$ of the blood diffusion model may or may not be practically identifiable at a single observational noise level.
@@ Appendix B Derivations of weak-form input-output equations for SIR model
$\begin{split}\int_{0}^{T}\ddot{\phi}I\textsf{d}t-\alpha^{2}\int_{0}^{T}\phi I\textsf{d}t&=\beta\int_{0}^{T}\dot{\phi}(\frac{I^{2}-(N-R(t)-I)^{2}}{2}\textsf{d}t\\ &-\beta\alpha\int_{0}^{T}\phi(I^{2}+2(N-R(t)-I)I)\textsf{d}t,\\ R(t)&=\int_{0}^{t}\alpha I(s)\textsf{d}s.\end{split}$
@@ Appendix E Comparison to Profile Likelihood
We compare the $(e,q)$-identifiability for the blood-diffusion model given by equation (7), found using the weak-form equation error method, to the profile likelihood (PLA), found using an output error method, for two different choices of parameters $w$. As can be seen in Figure 15, we $(10,q)$-identifiability of the blood diffusion model parameters and profile likelihood identifiability with $10\%$ observational error are in agreement. Further, for observational error under $10\%$ we find agreement between the WENDy confidence intervals and the PLA confidence intervals, which are given in Table 2. In a different region of the parameter space, we find the blood-diffusion model is not $(e,q)$-identifiable for observational error greater $0.5\%$ and that this practical non-identifiability is driven by uncertainty in parameter $w_{3}$. We find agreement for these results using the profile likelihood method at a $3\%$ observational error level. It is important to note that for the figures and confidence intervals generated using the profile likelihood method, the true parameters were used as an initial guess for the first output error optimization step in order to avoid failure of convergence, which was a challenge for the output error method.

### v5

@@ 3.2 Weak form Estimation of Nonlinear Dynamics (WENDy)
As presented in [BortzMessengerDukic2023BullMathBiol], the WENDy algorithm addresses the Errors-In-Variable nature of the linearized problem using iterative reweighting to approximate the correct underlying covariance structure [^8] The approximated covariance can also be used to calculate parameter confidence intervals. A brief derivation of the parameter covariance approximation is given in Appendix F, and we direct readers to [BortzMessengerDukic2023BullMathBiol] for further details.. This reweighting is dependent on the assumed noise structure of the observed data. For Gaussian additive noise, we apply the covariance correction using Algorithm 2 in [BortzMessengerDukic2023BullMathBiol]. Additionally, it has been shown in [RummelMessengerBeckerEtAl2025arXiv250208881] that, through a strategic choice of transformations to the observations $\mathbf{y}$, it is possible to approximate the covariance structure for systems where observational noise is multiplicative and lognormal (i.e. $\mathbf{y}=\mathbf{y}^{\star}\varepsilon$ where $\log(\varepsilon(\mathbf{t}))\sim\mathcal{N}(0,\sigma^{2})$).
@@ 3.4 Illustrating Examples
$\begin{split}\int_{0}^{T}\dot{\phi}I\textsf{d}t+\alpha\int_{0}^{T}\phi I\textsf{d}t&=-\beta\int_{0}^{T}\phi(R(t)+I-S_{0})I\textsf{d}t,\\ R(t)&=\int_{0}^{t}\alpha I(s)\textsf{d}s.\end{split}$ (17)
@@ 5 Conclusion
While global practical identifiability analysis is an emerging area of research [BholaDuraisamy2023SciRep, DobreBastogneRichard2010IFACProceedingsVolumes, ElWajehJungBongartzEtAl2022BullMathBiol] our $(e,q)$ criterion, like profile likelihood analysis and the FIM method, the $(e,q)$ method is local. This implies that, following best modeling practices, one may need to test a range of values in the parameter space rather than near a single parameterization of a system. For example, as we demonstrate in Appendix E, depending on the dynamics of the blood concentration, the parameter $w_{3}$ of the blood diffusion model may or may not be practically identifiable at a single observational noise level. However, we can also choose to assess $(e,q)$-identifiability on a range of parameters, rather than individual parameterizations of a specific model. In Figure 9, we show, using the blood-tissue diffusion model, for a range of parameter choices in $w_{2}$ and $w_{3}$, that if one parameterization yields an $(e,q)$-identifiable model, then a range of nearby parameterizations will also be $(e,q)$-identifiable. More generally, $(e,q)$-identifiability can be used in this way to explore a biologically informed parameter space.
Figure 9: The (e,q)-identifiability criterion holds locally in the parameter space. The minimum estimator error ratio, $q$, satisfied by the blood-tissue diffusion model for $w_{1}=30$ and varying parameters $w_{2}=[1,9]$ and $w_{3}=[10,20]$ using observational error ratios (A.) $e=1\%$, (B.) $e=5\%$, (C.) $e=10\%$, and (D.) $e=20\%$. The $(e,q)$-identifiability was determined from 1,000 simulations at each respective error level, $e$.
@@ Appendix B Derivations of weak-form input-output equations for SIR model
Derivation of Equation (17).
Consider the input-out representation of the SIR model given by equation (14). We divide by $I$ to obtain the following,
$0=\frac{\ddot{I}I-\dot{I}^{2}}{I^{2}}+\beta\alpha I+\beta\dot{I}.$ (18)
Notice that,
$\frac{d}{d\textsf{t}}(\frac{\dot{I}}{I})=\frac{\ddot{I}I-\dot{I}^{2}}{I^{2}}.$
Additionally, the equation (13) we know,
$\begin{split}\frac{dR}{d\textsf{t}}=\alpha I\Rightarrow R(t)=\int_{0}^{t}\alpha I(s)\textsf{d}s.\end{split}$
Note, here we assume a naive population, i.e., $R(0)=0$ and S(0)=NI(0)S(0)=N-I(0). Given that we observe the infected compartment, we are able to compute $R(t)$ through numeric integration. We are now able to generate the following equation from equation (18),
$\begin{split}\frac{d}{d\textsf{t}}(\frac{\dot{I}}{I})&=-\beta\frac{d}{d\textsf{t}}\left(\int_{0}^{t}\alpha I(s)d\textsf{s}+I\right)\\ &=-\beta\frac{d}{d\textsf{t}}(R(t)+I).\end{split}$ (19)
Note,
$\frac{\dot{I}}{I}=\frac{\beta SI-\alpha I}{I}=\beta S_{0}-\alpha.$
Next, we integrate equation (19) with respect to $t$ to obtain an alternative form of equation (14),
$\frac{\dot{I}}{I}=\beta(R(t)+I)+\beta S_{0}-\alpha.$
We then multiply by $I$,
$\dot{I}+\alpha I=-\beta(R(t)+I-S_{0})I.$ (20)
We obtain the weak-form input-output equation (17) by convoling equation (20) with a known test fucntion $\phi$.
An alternative form of Equation (17).
@@
Using WENDy with equation 30 to estimate parameter $\beta$ yields identical results to using WENDy with equation 17.
@@ Appendix E Comparison to Profile Likelihood
We compare the $(e,q)$-identifiability for the blood-diffusion model given by equation (7), found using the weak-form equation error method, to the profile likelihood (PLA), found using an output error method, for two different choices of parameters $w$. As can be seen in Figure 16, we $(10,q)$-identifiability of the blood diffusion model parameters and profile likelihood identifiability with $10\%$ observational error are in agreement. Further, for observational error under $10\%$ we find agreement between the WENDy confidence intervals and the PLA confidence intervals, which are given in Table 2. In a different region of the parameter space, we find the blood-diffusion model is not $(e,q)$-identifiable for observational error greater $0.5\%$ and that this practical non-identifiability is driven by uncertainty in parameter $w_{3}$. We find agreement for these results using the profile likelihood method at a $3\%$ observational error level. However, we note that a WENDy-based confidence interval is predicted for parameter $w_{3}$ despite both $(e,q)$ and PLA predicted non-identifiability. We hypothesize this discrepancy is driven by limitations in our approximation of the covariance matrix, which is described in Appendix F. It is important to note that for the figures and confidence intervals generated using the profile likelihood method, the true parameters were used as an initial guess for the first output error optimization step, to prevent the lsqnonlin optimization from failing to converge at higher levels of noise.
@@
Appendix F WENDy: Covariance Matrix
In this Appendix, we provide a brief description of the WENDy method approximation of the covariance matrix and derived parameter confidence intervals. A more detailed derivation of the covariance is provided in reference [BortzMessengerDukic2023BullMathBiol].
Consider the case of additive normal noise, i.e. $\mathbf{y}=\mathbf{y}^{*}+\varepsilon$ where $\varepsilon\sim\mathcal{N}(0,\sigma^{2}).$ The residual of equation $\mathbf{Gw^{*}=b}$ as defined in Section 3.2 is given below,
$\color[rgb]{0,0,0}\mathbf{r}(\mathbf{y,w})=\mathbf{G}(\mathbf{y})\mathbf{w}-\mathbf{b}(\mathbf{y}).$}
Both $\mathbf{G}$ and $\mathbf{b}$ act as nonlinear transformations of the noise present in the observational data $\mathbf{y}$. By taking the first-order Taylor expansion around the true output $\mathbf{y^{*}}$, we can obtain the following approximation of the residual,
$\color[rgb]{0,0,0}\mathbf{r}(\mathbf{y,w})\approx\mathbf{r}(\mathbf{y^{*},w})+J_{\mathbf{y}}(\mathbf{w})\varepsilon,$} (31)
where $J_{\mathbf{y}}$ is the Jacobian of $\mathbf{r}$ with respect to observational data. The covariance of the residuals is then approximated as
$\color[rgb]{0,0,0}\textbf{S}_{R}=\text{Cov}(\mathbf{r}(\mathbf{y,w}))\approx J_{\mathbf{y}}(\mathbf{w})\text{Cov}(\varepsilon)J_{\mathbf{y}}(\mathbf{w})^{T},$} (32)
where $\text{Cov}(\varepsilon)=\sigma^{2}\mathbf{I}_{M}.$
Using this covariance structure, WENDy applies an iteratively reweighted least squares algorithm to solve the following generalized least-squares problem
$\color[rgb]{0,0,0}\min_{\mathbf{w}}(\mathbf{Gw-b})^{T}\textbf{S}_{R}^{-1}(\mathbf{Gw-b}).$} (33)
From this relationship we can write the approximate parameter covariance matrix
$\color[rgb]{0,0,0}\textbf{S}_{w}\approx(\mathbf{G}^{T}\mathbf{G})^{-1}\mathbf{G}^{T}\textbf{S}_{R}\mathbf{G}(\mathbf{G}^{T}\mathbf{G})^{-1}.$} (34)
And, the $(1-\alpha)$ confidence interval around the parameter estimate $\hat{w}_{i}$
$\color[rgb]{0,0,0}[\hat{w}_{i}\pm F_{\textbf{S}_{w_{ii}}}(1-\alpha/2)],$} (35)
where $F$ is the CDF of the normal distribution with mean zero and variance $\textbf{S}_{w_{ii}}$.
In the case that we consider multiplicative lognormal noise, i.e. $\mathbf{y=y^{*}\varepsilon}$ where $\log(\varepsilon)\sim\mathcal{N}(0,\sigma^{2})$, we assume $\text{Cov}(\varepsilon)\approx\sigma^{2}\mathbf{diag}(y_{1},..,y_{M}),$ where $\mathbf{diag(x)}$ is a diagonal matrix with entries of $\mathbf{x}.$